OpenStudy (anonymous):

Choose the correct description of the graph of the inequality x > 3. Open circle on 3, shading to the left. Closed circle on 3, shading to the left. Open circle on 3, shading to the right. Closed circle on 3, shading to the right.

OpenStudy (compassionate):

The sign is grader than, so it's an open circle The open circle will be on 3 It will be shaded to the right.

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

help with another?

OpenStudy (compassionate):

I actually have to go! But ask @thomaster , he's good at this. :)

OpenStudy (anonymous):

@thomaster mind helping me??

thomaster (thomaster):

Lol srry I'm only good at algebra xD

OpenStudy (anonymous):

this is algebra? 0.o

OpenStudy (anonymous):

@amistre64 can you help me?

thomaster (thomaster):

lol serous? thought something with geometry xD

OpenStudy (anonymous):

nope its algebra 1

OpenStudy (anonymous):

Choose the correct description of the solution to the inequality -2x less than or greater to 10. Open circle on -5, shading to the left. Closed circle on -5, shading to the left. Open circle on -5, shading to the right. Closed circle on -5, shading to the right.

thomaster (thomaster):

Hm what do they mean with open circle?

OpenStudy (anonymous):

_-_ oh boy

thomaster (thomaster):

haha i'm not from the us you know :P

OpenStudy (kenljw):

Consider the real number line, x>3 indicates all numbers to the right of 3 in 2 coordinate system x^2 + y^2 = r^2 x= sqrt(r^2 -y^2) > 3 r^2 -y^2 > 9 if r = 3+ then y = 0 and you have the line The larger r gets the y limits are determine from above, not you never truly get a complete semi-circle. I assume open means not including and shading means not including.